\begin{thebibliography}{32}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Alaa and {van der Schaar}(2019)]{alaaDemystifyingBlackboxModels2019}
Ahmed~M. Alaa and Mihaela {van der Schaar}.
\newblock Demystifying {{Black-box Models}} with {{Symbolic Metamodels}}.
\newblock In \emph{Advances in {{Neural Information Processing Systems}}}, volume~32. {Curran Associates, Inc.}, 2019.

\bibitem[Angelis et~al.(2023)Angelis, Sofos, and Karakasidis]{angelisArtificialIntelligencePhysical2023}
Dimitrios Angelis, Filippos Sofos, and Theodoros~E. Karakasidis.
\newblock Artificial {{Intelligence}} in {{Physical Sciences}}: {{Symbolic Regression Trends}} and {{Perspectives}}.
\newblock \emph{Archives of Computational Methods in Engineering}, 30\penalty0 (6):\penalty0 3845--3865, July 2023.
\newblock ISSN 1886-1784.
\newblock \doi{10.1007/s11831-023-09922-z}.

\bibitem[{Bar-Sinai} et~al.(2019){Bar-Sinai}, Hoyer, Hickey, and Brenner]{bar-sinaiLearningDatadrivenDiscretizations2019}
Yohai {Bar-Sinai}, Stephan Hoyer, Jason Hickey, and Michael~P. Brenner.
\newblock Learning data-driven discretizations for partial differential equations.
\newblock \emph{Proceedings of the National Academy of Sciences}, 116\penalty0 (31):\penalty0 15344--15349, July 2019.
\newblock \doi{10.1073/pnas.1814058116}.

\bibitem[Battaglia et~al.(2018)Battaglia, Hamrick, Bapst, {Sanchez-Gonzalez}, Zambaldi, Malinowski, Tacchetti, Raposo, Santoro, and Faulkner]{battagliaRelationalInductiveBiases2018}
Peter~W. Battaglia, Jessica~B. Hamrick, Victor Bapst, Alvaro {Sanchez-Gonzalez}, Vinicius Zambaldi, Mateusz Malinowski, Andrea Tacchetti, David Raposo, Adam Santoro, and Ryan Faulkner.
\newblock Relational inductive biases, deep learning, and graph networks.
\newblock \emph{arXiv preprint arXiv:1806.01261}, 2018.

\bibitem[Brandstetter et~al.(2022)Brandstetter, Worrall, and Welling]{brandstetterMessagePassingNeural2022}
Johannes Brandstetter, Daniel~E. Worrall, and Max Welling.
\newblock Message {{Passing Neural PDE Solvers}}.
\newblock In \emph{International {{Conference}} on {{Learning Representations}}}, January 2022.

\bibitem[Chrysos et~al.(2021)Chrysos, Moschoglou, Bouritsas, Deng, Panagakis, and Zafeiriou]{chrysosDeepPolynomialNeural2021}
Grigorios Chrysos, Stylianos Moschoglou, Giorgos Bouritsas, Jiankang Deng, Yannis Panagakis, and Stefanos Zafeiriou.
\newblock Deep {{Polynomial Neural Networks}}.
\newblock \emph{IEEE Transactions on Pattern Analysis and Machine Intelligence}, pages 1--1, 2021.
\newblock ISSN 0162-8828, 2160-9292, 1939-3539.
\newblock \doi{10.1109/TPAMI.2021.3058891}.

\bibitem[Cranmer et~al.(2019)Cranmer, Xu, Battaglia, and Ho]{cranmerLearningSymbolicPhysics2019}
Miles~D. Cranmer, Rui Xu, Peter Battaglia, and Shirley Ho.
\newblock Learning symbolic physics with graph networks.
\newblock \emph{arXiv preprint arXiv:1909.05862}, 2019.

\bibitem[Cuomo et~al.(2022)Cuomo, {di Cola}, Giampaolo, Rozza, Raissi, and Piccialli]{cuomoScientificMachineLearning2022}
Salvatore Cuomo, Vincenzo~Schiano {di Cola}, Fabio Giampaolo, Gianluigi Rozza, Maziar Raissi, and Francesco Piccialli.
\newblock Scientific {{Machine Learning}} through {{Physics-Informed Neural Networks}}: {{Where}} we are and {{What}}'s next, June 2022.

\bibitem[Dissanayake and {Phan-Thien}(1994)]{dissanayakeNeuralnetworkbasedApproximationsSolving1994}
{\relax MWMG}~Dissanayake and Nhan {Phan-Thien}.
\newblock Neural-network-based approximations for solving partial differential equations.
\newblock \emph{Communications in Numerical Methods in Engineering}, 10\penalty0 (3):\penalty0 195--201, 1994.

\bibitem[Fanaskov and Oseledets(2022)]{fanaskovSpectralNeuralOperators2022}
V.~Fanaskov and I.~Oseledets.
\newblock Spectral {{Neural Operators}}, May 2022.

\bibitem[Fronk and Petzold(2023)]{fronkInterpretablePolynomialNeural2023}
Colby Fronk and Linda Petzold.
\newblock Interpretable {{Polynomial Neural Ordinary Differential Equations}}.
\newblock \emph{Chaos: An Interdisciplinary Journal of Nonlinear Science}, 33\penalty0 (4):\penalty0 043101, April 2023.
\newblock ISSN 1054-1500, 1089-7682.
\newblock \doi{10.1063/5.0130803}.

\bibitem[Goswami et~al.(2022)Goswami, Bora, Yu, and Karniadakis]{goswamiPhysicsInformedDeepNeural2022}
Somdatta Goswami, Aniruddha Bora, Yue Yu, and George~Em Karniadakis.
\newblock Physics-{{Informed Deep Neural Operator Networks}}, July 2022.

\bibitem[Hsieh et~al.(2019)Hsieh, Zhao, Eismann, Mirabella, and Ermon]{hsiehLearningNeuralPDE2019}
Jun-Ting Hsieh, Shengjia Zhao, Stephan Eismann, Lucia Mirabella, and Stefano Ermon.
\newblock Learning {{Neural PDE Solvers}} with {{Convergence Guarantees}}.
\newblock In \emph{International {{Conference}} on {{Learning Representations}}}, April 2019.

\bibitem[Koza(1992)]{KozaProgrammingComputers}
JRGP Koza.
\newblock On the programming of computers by means of natural selection.
\newblock \emph{Genetic programming}, 1992.

\bibitem[Krishnapriyan et~al.(2021)Krishnapriyan, Gholami, Zhe, Kirby, and Mahoney]{krishnapriyanCharacterizingPossibleFailure2021}
Aditi~S. Krishnapriyan, Amir Gholami, Shandian Zhe, Robert~M. Kirby, and Michael~W. Mahoney.
\newblock Characterizing possible failure modes in physics-informed neural networks, November 2021.

\bibitem[Lagaris et~al.(Sept./1998)Lagaris, Likas, and Fotiadis]{lagarisArtificialNeuralNetworks1998}
I.~E. Lagaris, A.~Likas, and D.~I. Fotiadis.
\newblock Artificial {{Neural Networks}} for {{Solving Ordinary}} and {{Partial Differential Equations}}.
\newblock \emph{IEEE Transactions on Neural Networks}, 9\penalty0 (5):\penalty0 987--1000, Sept./1998.
\newblock ISSN 10459227.
\newblock \doi{10.1109/72.712178}.

\bibitem[Lee et~al.(2021)Lee, Trask, and Stinis]{leeStructurepreservingSparseIdentification2021b}
Kookjin Lee, Nathaniel Trask, and Panos Stinis.
\newblock Structure-preserving {{Sparse Identification}} of {{Nonlinear Dynamics}} for {{Data-driven Modeling}}, September 2021.

\bibitem[Li et~al.(2020)Li, Kovachki, Azizzadenesheli, Liu, Bhattacharya, Stuart, and Anandkumar]{liNeuralOperatorGraph2020}
Zongyi Li, Nikola Kovachki, Kamyar Azizzadenesheli, Burigede Liu, Kaushik Bhattacharya, Andrew Stuart, and Anima Anandkumar.
\newblock Neural {{Operator}}: {{Graph Kernel Network}} for {{Partial Differential Equations}}, March 2020.

\bibitem[Li et~al.(2021{\natexlab{a}})Li, Kovachki, Azizzadenesheli, Liu, Bhattacharya, Stuart, and Anandkumar]{liFourierNeuralOperator2021}
Zongyi Li, Nikola Kovachki, Kamyar Azizzadenesheli, Burigede Liu, Kaushik Bhattacharya, Andrew Stuart, and Anima Anandkumar.
\newblock Fourier {{Neural Operator}} for {{Parametric Partial Differential Equations}}, May 2021{\natexlab{a}}.

\bibitem[Li et~al.(2021{\natexlab{b}})Li, Zheng, Kovachki, Jin, Chen, Liu, Azizzadenesheli, and Anandkumar]{liPhysicsinformedNeuralOperator2021}
Zongyi Li, Hongkai Zheng, Nikola Kovachki, David Jin, Haoxuan Chen, Burigede Liu, Kamyar Azizzadenesheli, and Anima Anandkumar.
\newblock Physics-informed neural operator for learning partial differential equations.
\newblock \emph{arXiv preprint arXiv:2111.03794}, 2021{\natexlab{b}}.

\bibitem[Lu et~al.(2021)Lu, Jin, and Karniadakis]{luDeepONetLearningNonlinear2021}
Lu~Lu, Pengzhan Jin, and George~Em Karniadakis.
\newblock {{DeepONet}}: {{Learning}} nonlinear operators for identifying differential equations based on the universal approximation theorem of operators.
\newblock \emph{Nature Machine Intelligence}, 3\penalty0 (3):\penalty0 218--229, March 2021.
\newblock ISSN 2522-5839.
\newblock \doi{10.1038/s42256-021-00302-5}.

\bibitem[Lu et~al.(2022)Lu, Meng, Cai, Mao, Goswami, Zhang, and Karniadakis]{luComprehensiveFairComparison2022}
Lu~Lu, Xuhui Meng, Shengze Cai, Zhiping Mao, Somdatta Goswami, Zhongqiang Zhang, and George~Em Karniadakis.
\newblock A comprehensive and fair comparison of two neural operators (with practical extensions) based on fair data.
\newblock \emph{Computer Methods in Applied Mechanics and Engineering}, 393:\penalty0 114778, 2022.

\bibitem[Meng et~al.(2022)Meng, Seo, Cao, Griesemer, and Liu]{mengWhenPhysicsMeets2022}
Chuizheng Meng, Sungyong Seo, Defu Cao, Sam Griesemer, and Yan Liu.
\newblock When {{Physics Meets Machine Learning}}: {{A Survey}} of {{Physics-Informed Machine Learning}}, March 2022.

\bibitem[Meuris et~al.(2023)Meuris, Qadeer, and Stinis]{meurisMachinelearningbasedSpectralMethods2023}
Brek Meuris, Saad Qadeer, and Panos Stinis.
\newblock Machine-learning-based spectral methods for partial differential equations.
\newblock \emph{Scientific Reports}, 13\penalty0 (1):\penalty0 1739, January 2023.
\newblock ISSN 2045-2322.
\newblock \doi{10.1038/s41598-022-26602-3}.

\bibitem[Pathak et~al.(2022)Pathak, Subramanian, Harrington, Raja, Chattopadhyay, Mardani, Kurth, Hall, Li, Azizzadenesheli, Hassanzadeh, Kashinath, and Anandkumar]{pathakFourCastNetGlobalDatadriven2022}
Jaideep Pathak, Shashank Subramanian, Peter Harrington, Sanjeev Raja, Ashesh Chattopadhyay, Morteza Mardani, Thorsten Kurth, David Hall, Zongyi Li, Kamyar Azizzadenesheli, Pedram Hassanzadeh, Karthik Kashinath, and Animashree Anandkumar.
\newblock {{FourCastNet}}: {{A Global Data-driven High-resolution Weather Model}} using {{Adaptive Fourier Neural Operators}}, February 2022.

\bibitem[Psichogios and Ungar(1992)]{psichogiosHybridNeuralNetworkfirst1992}
Dimitris~C. Psichogios and Lyle~H. Ungar.
\newblock A hybrid neural network-first principles approach to process modeling.
\newblock \emph{AIChE Journal}, 38\penalty0 (10):\penalty0 1499--1511, 1992.

\bibitem[Rahman et~al.(2022)Rahman, Ross, and Azizzadenesheli]{rahmanUshapedNeuralOperators2022}
Md~Ashiqur Rahman, Zachary~E. Ross, and Kamyar Azizzadenesheli.
\newblock U-{{NO}}: {{U-shaped Neural Operators}}, May 2022.

\bibitem[Raissi et~al.(2019)Raissi, Perdikaris, and Karniadakis]{raissiPhysicsinformedNeuralNetworks2019}
M.~Raissi, P.~Perdikaris, and G.~E. Karniadakis.
\newblock Physics-informed neural networks: {{A}} deep learning framework for solving forward and inverse problems involving nonlinear partial differential equations.
\newblock \emph{Journal of Computational Physics}, 378:\penalty0 686--707, February 2019.
\newblock ISSN 0021-9991.
\newblock \doi{10.1016/j.jcp.2018.10.045}.

\bibitem[{Sanchez-Gonzalez} et~al.(2018){Sanchez-Gonzalez}, Heess, Springenberg, Merel, Riedmiller, Hadsell, and Battaglia]{sanchez-gonzalezGraphNetworksLearnable2018}
Alvaro {Sanchez-Gonzalez}, Nicolas Heess, Jost~Tobias Springenberg, Josh Merel, Martin Riedmiller, Raia Hadsell, and Peter Battaglia.
\newblock Graph networks as learnable physics engines for inference and control.
\newblock In \emph{International {{Conference}} on {{Machine Learning}}}, pages 4470--4479. {PMLR}, 2018.

\bibitem[Thuerey et~al.(2022)Thuerey, Holl, Mueller, Schnell, Trost, and Um]{thuereyPhysicsbasedDeepLearning2022}
Nils Thuerey, Philipp Holl, Maximilian Mueller, Patrick Schnell, Felix Trost, and Kiwon Um.
\newblock Physics-based {{Deep Learning}}, April 2022.

\bibitem[Tripura and Chakraborty(2022)]{tripuraWaveletNeuralOperator2022}
Tapas Tripura and Souvik Chakraborty.
\newblock Wavelet neural operator: A neural operator for parametric partial differential equations, May 2022.

\bibitem[Um et~al.(2020)Um, Brand, Fei, Holl, and Thuerey]{umSolverintheLoopLearningDifferentiable2020}
Kiwon Um, Robert Brand, Yun~(Raymond) Fei, Philipp Holl, and Nils Thuerey.
\newblock Solver-in-the-{{Loop}}: {{Learning}} from {{Differentiable Physics}} to {{Interact}} with {{Iterative PDE-Solvers}}.
\newblock In \emph{Advances in {{Neural Information Processing Systems}}}, volume~33, pages 6111--6122. {Curran Associates, Inc.}, 2020.

\end{thebibliography}

\documentclass{article}

\PassOptionsToPackage{numbers, sort&compress}{natbib}
\usepackage[final]{neurips_2021}

\usepackage[utf8]{inputenc} 
\usepackage[T1]{fontenc}    
\usepackage{hyperref}       
\usepackage{url}            
\usepackage{booktabs}       
\usepackage{amsfonts}       
\usepackage{nicefrac}       
\usepackage{microtype}      
\usepackage{xcolor}         
\usepackage{glossaries}

\newacronym{pinn}{PINN}{physics-informed neural network}
\newacronym{mpinn}{M-PINN}{manifold PINN}
\newacronym{cpinn}{C-PINN}{conformal PINN}
\newacronym{mppde}{MP-PDE}{message passing neural PDE solver}
\newacronym{fno}{FNO}{Fourier neural operator}
\newacronym{pino}{PINO}{physics-informed neural operator}
\newacronym{wno}{WNO}{wavelet neural operator}
\newacronym{dsn}{DSN}{differential spectral normalization}
\newacronym{sindy}{SINDy}{sparse identification of nonlinear dynamics}
\newacronym{siren}{SIREN}{sinusoidal representation network}
\newacronym{mlp}{MLP}{multilayer perceptron}
\newacronym{cnn}{CNN}{convolutional neural network}
\newacronym{rnn}{RNN}{recurrent neural network}
\newacronym{pca}{PCA}{principal component analysis}
\newacronym{fft}{FFT}{fast Fourier transform}
\newacronym{inr}{INR}{implicit neural representation}
\newacronym{ginr}{GINR}{generalised implicit neural representation}
\newacronym{pde}{PDE}{partial differential equation}
\newacronym{ode}{ODE}{ordinary differential equation}
\newacronym{resnet}{ResNet}{residual network}
\newacronym{ml}{ML}{machine learning}
\newacronym{grand}{GRAND}{graph neural diffusion}
\newacronym{gcn}{GCN}{graph convolutional network}
\newacronym{gcnn}{G-CNN}{group-equivariant convolutional neural networks}
\newacronym{gnn}{GNN}{graph neural network}

\title{Interpretable Neural PDE Solvers using Symbolic Frameworks}

\author{%
  Yolanne Yi Ran Lee\\
  Department of Computer Science\\
  University College London\\
  Gower Street, London WC1E 6BT, UK \\
  \texttt{yolanne.lee.19@ucl.ac.uk}
}

\begin{document}

\maketitle

\begin{abstract}
  Partial differential equations (PDEs) are ubiquitous in the world around us, modelling phenomena from heat and sound to quantum systems. Recent advances in deep learning have resulted in the development of powerful neural solvers; however, while these methods have demonstrated state-of-the-art performance in both accuracy and computational efficiency, a significant challenge remains in their interpretability. Most existing methodologies prioritize predictive accuracy over clarity in the underlying mechanisms driving the model's decisions. Interpretability is crucial for trustworthiness and broader applicability, especially in scientific and engineering domains where neural PDE solvers might see the most impact. In this context, a notable gap in current research is the integration of symbolic frameworks (such as symbolic regression) into these solvers. Symbolic frameworks have the potential to distill complex neural operations into human-readable mathematical expressions, bridging the divide between black-box predictions and solutions.
\end{abstract}

\section{Introduction}

\Glspl{pde} are ubiquitous to modern physics and engineering. Though they have existed since the first system of equations found by Euler over 250 years ago, many challenges remain in finding numerical solutions resolving spatiotemporal features over multiple scales and nonlinear stochastic systems. Currently, numerical solvers such as finite differences and finite elements can solve \glspl{pde} when analytical solutions cannot be found, like in nonlinear and high dimensional systems, by discretizing a problem over a grid and evolving over time at very fine-grain timesteps. However, such solvers are computationally complex even with modern computers, and can fail when faced by multiscale and/or stochastic \glspl{pde}.

\subsection{Neural PDE Solvers}

Neural networks can learn and generalize to new contexts such as different initial/boundary conditions, coefficients, or even different \glspl{pde} entirely. They are a particularly promising avenue for solving highly complex \glspl{pde} such as those found in weather prediction and fluid dynamics. For a review which contextualizes physics informed machine learning with regards to classical problems and methods, see \cite{mengWhenPhysicsMeets2022}.

Today, neural PDE solvers are capable of remarkable performance in historically challenging tasks. FourCastNet, proposed by \citet{pathakFourCastNetGlobalDatadriven2022}, achieves comparable short-term forecasting precision to the ECMWF Integrated Forecasting System (IFS), a cutting-edge numerical weather prediction (NWP) model, for large-scale variables like atmospheric pressure. It surpasses IFS in predicting small-scale variables, such as precipitation. Most notably, however, is the fact that it is 4 to 5 orders of magnitude faster than most NWP models and uses a fraction of the number of variables to achieve its remarkable performance. This success is based off of the popular neural operator framework and, more specifically, \glspl{fno} \cite{liFourierNeuralOperator2021}.

However, models designed to solve more fundamental mathematical tasks like \glspl{pde} directly do not see as much acclaim or practical adoption. While we use \glspl{pde} to model the mathematics underlying heat, sound, electrodynamics, fluid dynamics, and many other physical phenomena which are essential to many engineering and medical tasks (among other fields), one major issue hinders real-world applications of neural methods: interpretability.

Neural \gls{pde} solvers lack the guarantees and transparency that numerical solvers have. This prevents them to be used in what are often high-stakes applications like engineering and medical simulations, even despite any potential gains in accuracy, generalisability, or computation time.

Now that we have demonstrated the clear potential for cutting edge neural \gls{pde} solvers, some of which are laid out below, focus should move toward bringing these methods closer to real-world application by considering needs beginning with interpretability and extending to more stringent demands such as performance bounds or limits.

\section{An overview of the current methods}

The idea of using deep learning techniques to solve differential equations has a long history, including Dissanayake's and Phan-Thien's attempt to use \glspl{mlp} as universal approximators to solve \glspl{pde}, and arguably includes any work involving incorporating prior knowledge into models in general \cite{dissanayakeNeuralnetworkbasedApproximationsSolving1994, psichogiosHybridNeuralNetworkfirst1992, lagarisArtificialNeuralNetworks1998}. Although some hybrid classical-neural methods exist \cite{hsiehLearningNeuralPDE2019,umSolverintheLoopLearningDifferentiable2020,meurisMachinelearningbasedSpectralMethods2023,bar-sinaiLearningDatadrivenDiscretizations2019}, which are inherently interpretable to varying degrees, we focus on fully neural methods: methods which rely on the universal function approximation theory, such that a sufficiently complex network can represent any arbitrary function. One very popular method mentioned already above is the neural operator, which models the solution of a \gls{pde} as an operator mapping inputs to outputs. The problem is set such that a neural operator $\mathcal{M}$ satisfies $\mathcal{M}(t,\mathbf{u}^{0}) = \mathbf{u}(t)$ where $\mathbf{u}$ is the solution, and $\mathbf{u}^{0}$ are the initial conditions \cite{luDeepONetLearningNonlinear2021, brandstetterMessagePassingNeural2022}. Simple \glspl{mlp}, \glspl{cnn}, \glspl{rnn}, and other networks used to map input vectors to output vectors are naive examples of finite-dimensional operators.

\subsection{Physics Informed Neural Networks}

In 2017, Raissi et al. introduced the \gls{pinn} \cite{raissiPhysicsinformedNeuralNetworks2019}. They structure the problem such that the network, denoted as \( \mathcal{N} \), fulfills \( \mathcal{N}(t, \mathbf{u}^{0}) = \mathbf{u}(t) \), with \( \mathbf{u}^{0} \) representing the initial conditions. The core idea of \glspl{pinn} is to directly integrate relevant physical laws into the network’s predictions, which is achieved by incorporating additional loss term(s) into the network's objective function.


For a typical loss function
$\theta = \text{argmin}_{\theta} \mathcal{L}(\theta)$

the loss with a physics prior may be defined as follows:

\begin{equation} \label{eq:pinnloss}
    \mathcal{L}(\theta) = \omega_{\mathcal{F}} \mathcal{L}_{\mathcal{F}}(\theta) + \omega_{\mathcal{B}} \mathcal{L}_{\mathcal{B}}(\theta) + \omega_{d} \mathcal{L}_{\text{data}}(\theta)
\end{equation}

where $\mathcal{L}_{\mathcal{B}}$ penalises against the initial and/or boundary conditions to fit the known data over the network, $\mathcal{L}_{\mathcal{F}}$ enforces the \gls{pde} itself at collocation points (which are calculated using auto-differentiation to compute derivatives of $\mathbf{\hat{u}_{\theta}(\mathbf{z})}$), and $\mathcal{L}_{\text{data}}$ (the standard loss) forces $\mathbf{\hat{u}}_{\theta}$ to match measurements of $\mathbf{u}$ over the provided data points.

Since the network maps input variables to output variables which are both finite-dimensional and dependent on the grid used to discretize the problem domain, it is considered a finite dimensional neural operator. The paper gained a lot of traction and inspired many architectures which now fall under the \gls{pinn} family; for a more thorough review, see \cite{cuomoScientificMachineLearning2022}, and for hands-on examples visit this digital book \cite{thuereyPhysicsbasedDeepLearning2022}.

The success of this loss-based approach is apparent when considering the rapid growth of papers which extend the original iteration of the \gls{pinn}. It is conceptually interpretable (though its performance pales in comparison to later methods) and can be simple to implement.

However, Krishnapriyan et al. \cite{krishnapriyanCharacterizingPossibleFailure2021} has shown that even though standard fully-connected neural networks are theoretically capable of representing any function given enough neurons and layers, a \gls{pinn} may still fail to approximate a solution due to the complex loss landscapes arising from soft \gls{pde} constraints.

\subsection{DeepONets}

The DeepONet architecture is a seminal example of an infinite dimensional neural operator in contrast to the finite dimensional \gls{pinn} \cite{luDeepONetLearningNonlinear2021}. It consists of one or multiple branch nets which encode discrete inputs to an input function space, and a single trunk net which receives the query location to evaluate the output function. The model maps from a fixed, finite dimensional grid to an infinite dimensional output space.

Since the development of the DeepONet, many novel neural operators have emerged which generalize this finite-infinite dimensional mapping to an infinite-infinite dimensional mapping \cite{liNeuralOperatorGraph2020, liPhysicsinformedNeuralOperator2021, goswamiPhysicsInformedDeepNeural2022, rahmanUshapedNeuralOperators2022, tripuraWaveletNeuralOperator2022, fanaskovSpectralNeuralOperators2022, pathakFourCastNetGlobalDatadriven2022}, including the \gls{fno} \cite{liFourierNeuralOperator2021}. This network operates within Fourier space and takes advantage of the convolution theorem to place the integral kernel in Fourier space as a convolutional operator.

\subsection{Fourier Neural Operators}

Concretely, the mapping between two infinite-dimensional spaces is learned from the discrete number of observed pairs. These global integral operators (implemented as Fourier space convolutional operators) are combined with local nonlinear activation functions, resulting in an architecture which is highly expressive yet computationally efficient, as well as being resolution-invariant. The \gls{fno} acts as a DeepONet with the branch net approximating the input functions and the trunk net using Fourier basis functions.

While the vanilla \gls{fno} required the input function to be defined on a grid due to its reliance on the \gls{fft}, further work developed mesh-independent variations as well \cite{liPhysicsinformedNeuralOperator2021}. In brief, the d\gls{fno}+ and g\gls{fno}+ presented in \cite{luComprehensiveFairComparison2022} which adapt the \gls{fno} to nonlinear mappings and complex geometries respectively. \citet{tripuraWaveletNeuralOperator2022} present the \gls{wno} which learns the network parameters in wavelet space rather than Fourier space to allow for both frequency and spatial resolution (the latter of which Fourier basis functions are not explicitly able to capture). The \gls{wno} was shown to handle highly nonlinear \glspl{pde} with sharp changes and discontinuities on both smooth and complex geometries. The \gls{pino} \cite{liPhysicsinformedNeuralOperator2021}, combines the \gls{pinn} methodology with the \gls{fno} resulting in an architecture which ameliorates the optimization challenges that \glspl{pinn} face, as well as improving the accuracy of both models. In short, the \gls{pino} first learns a solution operator using some combination of the data loss and/or \gls{pde} loss, and then uses the learned operator as an ansatz to approximate the ground truth operator with the \gls{pde} loss (and, optionally, the operator loss from the previous step).

Neural methods are often able to operate on multiple domains, computationally efficient at test time, and can be completely data-driven. However, none of the above popular examples have a clear interpretation.

\section{Symbolic Programming}

Symbolic regression was first proposed by \citet{KozaProgrammingComputers} and is synonymous to regular linear or nonlinear regression except it fits over the parameters and structure of an equation rather than numerical data points. An equation can be represented as an acyclic graph, with its leaves being the operations and its terminal nodes being variables and constants. The method generates random operations as the base structure of the graph, and uses genetic programming techniques such as point mutation and crossover to introduce variation (exploration of the search space) \cite{angelisArtificialIntelligencePhysical2023}. Each equation is assessed against how many data points it can handle, often with an error metric defined using mean squared error or other standard forms. As promising equations are iterated and evaluated, the algorithm converges toward the resulting symbolic equation that best describes the data. For a more detailed overview of symbolic regression, \citet{angelisArtificialIntelligencePhysical2023} presents a recent overview including the applications in machine learning to date.

Numerical models have traditionally been developed systematically by hand and with a deep understanding of the system and its setting. Accordingly, fewer models exist for larger systems or those that are less explored. With the still-growing amount of compute available now, there is new potential for data-driven methods to discover relationships in data. Neural networks have been combined with symbolic regression in a variety of ways.

\citet{cranmerLearningSymbolicPhysics2019} capitalize on the development of \glspl{gnn} \cite{battagliaRelationalInductiveBiases2018, sanchez-gonzalezGraphNetworksLearnable2018} such that the \gls{gnn} has some inductive biases imposed to mimic some true latent space. This model is then trained to predict the dynamics of common physical systems, and symbolic regression is finally used to extract algebraic equations from the \gls{gnn} messages and inputs. As an example, the update to some physical body's velocity may be calculated using Euler integration

\begin{equation}
    \mathbf{v}_{i}^{'} = \mathbf{v}_{i} + \phi^{v}(\mathbf{v}_{i},\mathbf{\bar{e}_{i}^{'}})
\end{equation}

where $\phi^{v}$ is the node update function which takes the node updates $\mathbf{v}_{i}$ and the pooled messages $\mathbf{\bar{e}_{i}^{'}}$ at the $i$-th receiver node. For the ideal case that the \gls{gnn} predicts $\mathbf{v}_{i}^{'}$, the node update over all received and pooled messages must then be equal to the net force acting on the body. The message vectors therefore would be linear transformations of the forces, which can be extracted as symbolic expressions using symbolic regression.

\section{The Potential for Symbolic Frameworks in Neural PDE Solvers}

Some work in this direction has begun to emerge. \citet{leeStructurepreservingSparseIdentification2021b} interprets the neural ODE by using its resulting time derivative predictions as inputs to SINDy (a framework based on sparse regression to discover parsimonious governing equations including, but not limited to \glspl{pde}) which interprets the results as a symbolic equation. In this case, two outputs are expected: the solutions across however many time steps the network predicts, and also a symbolic equation describing the former. Another similar approach by \citet{fronkInterpretablePolynomialNeural2023} adapts the Deep Polynomial Neural Networks from \cite{chrysosDeepPolynomialNeural2021}, which constrains the output vector such that each element is represented as a polynomial based on every input element, directly for symbolic regression. Then, this network is embedded into the neural ODE to design a network capable of learning polynomial differential equations, such as those found in certain dynamical systems. 

While the former two approaches add or otherwise modify an existing architecture, \cite{alaaDemystifyingBlackboxModels2019} presents a "model-of-a-model" approach called the symbolic metamodel, which takes a learned model as input and outputs a symbolic equation as a post-hoc interpretation. The symbolic metamodel is based on the Meijer G-function, $G_{m,n}^{p,q} (a_p, b_q | x)$, where $\mathbf{a}$ and $\mathbf{b}$ are two sets of real-values which specify a particular instance of $G$. One limitation is that this method cannot represent the outputs by means of any differential equations due to the limitations of the Meijer G-function. Since the target symbolic equation can be expressed (within this limited scope) as a Meijer G-function, it can be learned using standard gradient descent methods instead of symbolic regression.

However, further integrations would greatly benefit popular models like the neural operator methods described above, among others. In fact, closer interdisciplinary work involving "XAI" (explainable AI) methodology will also have bidirectional benefit. Neural solvers such as these which tackle rigorous problems such as complex \glspl{pde} pose a great stepping stone problem since it uniquely presents many ways to evaluate the correctness of an interpretation and its fitness to the network itself. Being able to verify interpretation techniques in this type of problem also makes the techniques more trustworthy to be used in less straightforward problem settings more generally.

\subsection{Considerations for further study}

Key performance indicators, such as predictive accuracy, computational efficiency, and convergence speed, need to be thoroughly assessed to determine the effectiveness and reliability of the integrated model. As in the standard neural \gls{pde} solvers, striking a balance between high-accuracy predictions and computational efficiency. Furthermore, exploring how well these models generalize across varied problem domains, scales, and complexities would offer insights into their robustness and versatility, thus making them suitable for a broader spectrum of practical applications.

Addressing the interpretability and transparency of the model's predictions stands as another cornerstone in this integration. Embedding symbolic frameworks should not only facilitate accurate and computationally efficient PDE solutions but more importantly provide clear, understandable mathematical formulations that highlight the underlying relationships and dynamics captured by the neural network. As such, the symbolic expressions derived from the model should be mathematically coherent, ideally providing a lens through which the model’s predictions and decision-making processes can be analyzed and validated against known physical laws and empirical observations.

\section{Summary}

Enhanced interpretability is crucial, especially in science and engineering, where understanding the underlying processes and making sure our predictions are reliable is vital. By integrating symbolic programming with neural \gls{pde} solvers, we ensure that the models we create are not just data-driven black boxes. Instead, they have enhanced visibility thanks to their ability to respect and show physical laws and principles, filling the gap between data-driven machine learning and scientific modelling guided by theories. So, combining symbolic frameworks and neural \gls{pde} solvers opens up a promising path. It brings together the strength of machine learning and the clarity and accuracy of symbolic math expressions, pushing forward our abilities in scientific computing and simulation.

\bibliography{bib}

\end{document}